# Path-integral Monte Carlo estimator for the dipole polarizability of quantum plasma

### v1

@@ IV Results
To make physical benchmarks, there is obvious systematic bias related to the quantum statistics, i.e., the lack of Fermion exchange in the simulation. Otherwise, the Coulomb pair-density matrix is numerically exact and there are no nodal constraints. Remaining bias ensues from using a finite time-step $\Delta\tau$ to treat higher-body correlations (beyond pair approximation) and periodic images. As the higher-body correlations intensify with density, we assume for this work a choice of the time-step according to $\Delta\tau=ars^{3}$, where $a=0.001$ Ha-1.
@@
The finite-size effects of the dipole polarizability are more topical: The scaling of $\tilde{G}_{\mu\mu}(\tau)$ and $\tilde{G}_{\mu\mu}(i\omega_{n})$ versus $N$ is plotted in Fig. 3 for $k_{c}r_{c}=15.0$. Overall, the effects of $\tilde{G}_{\mu\mu}$ from PIMC versus Drude due to finite size are negligible regardless of $N$. Nevertheless, we will use $N=16$ for the remainder of this work. The effects due to finite time-step and the Ewald cutoff parameter $k_{c}r_{c}$ are also statistically insignificant.
The lack of finite-size effects in $\tilde{G}_{\mu\mu}$ is in stark contrast with the energies, which are affected by the correlation and, although absent here, exchange. The indifference to the number of particles is therefore a manifestation of a greater principle: the trajectories of an ensemble of free particles in a periodic setting sample the phase space in collective harmony. The dipole autocorrelation estimator only measures the collective state and distributes it evenly for each particle. Another way of seeing this is that the low-momentum states of the plasma are occupied and cannot be used to excite the dielectric response of individual particles. Curiously, the cases of interacting (PIMC) and non-interacting (Drude) come out effectively the same.
@@
Figure 4: On the left, $\tilde{G}_{\mu\mu}(\tau)$ and $\Delta\tilde{G}_{\mu\mu}(\tau)$ with variable $\Theta$ at $r_{s}=4$. On the right, $\tilde{G}_{\mu\mu}(i\omega_{n})$ and $\Delta\tilde{G}_{\mu\mu}(i\omega_{n})$ with variable $\Theta$ at $r_{s}=4$. Solid lines represent the absolute data and dotted lines the statistical uncertainties.
Finally, we study how $\tilde{G}_{\mu\mu}(\tau)$ depends on physical parameters, the finite temperature and density. In Fig. 4 we plot $\tilde{G}_{\mu\mu}(\tau)$ and $\tilde{G}_{\mu\mu}(i\omega_{n})$ versus $\Theta$ with $r_{s}=4$ and $N=16$. The plotting axes of $\tilde{G}_{\mu\mu}(\tau)$ are rescaled to $\tau/\beta$, emphasizing that the absolute values of $\tilde{G}_{\mu\mu}(\tau=\beta/2)$ decrease with the inverse temperature. Consequently, the Matsubara integral $\tilde{G}(i\omega_{n}=0)$ goes down with $\beta$, but otherwise the Matsubara series $\tilde{G}(i\omega_{n}>0)$ follow the same curve, only with different spacings. Clearly, the PIMC results match the Drude model within statistical uncertainty of $2\%$ or less.
In Figure 5 we discuss how the match depends on the Drude relaxation parameter $\Gamma$. When $\Gamma$ is increased, the Drude correlation function $\tilde{G}_{\mu\mu}(\tau)$ loses magnitude. The phenomenological explanation is that the scattering rate of the charge carriers from ions increases with $\Gamma$, and thus, the spatial diffusion slows down. The PIMC simulation of free particles has no scattering, and thus, the correspondence is the best with low $\Gamma$. This is presented in the lower panel of Fig. 5, where linear fits approach $\eta\rightarrow 0$ when $\Gamma\rightarrow 0$. The relative error $\eta$ scales linearly with $\Gamma$ and $\beta$, and so, curves representing $\eta/\beta$ are overlaid within statistical uncertainties. With this consideration we want to suggest a future possibility of estimating $\Gamma$ from PIMC simulations, provided a reasonable but yet unknown sampling method of the scattering process. As seen from all the presented results, such a simulation could be easily validated, while providing first-principles insights to the Drude scattering. Furthermore, using a finite-$\Gamma$ simulation such as that could also lend itself for studying nonlinear optical phenomena of Drude-like materials on a quantum mechanical level.
@@
Figure 5: On the left, a comparison of $\tilde{G}_{\mu\mu}(\tau)$ versus $\tau$ from PIMC (solid black line) and the Drude model with variable values for $\Gamma$. On the right, the rescaled goodness measure $\eta/\beta$ is plotted against $\Gamma$ based on variables temperatures and $r_{s}=4$. Dotted lines are linear fits to the data.
@@ V Summary and outlook
We have investigated a PIMC estimator for the dipole polarizability of periodic quantum gases and plasmas in the long-wavelength limit, i.e., the optical regime. We have bench marked the estimator with quantum Coulomb plasma in a range of finite densities ($r_{s}=2\ldots 8$ a0) and finite temperatures ($\Theta=0.1\ldots 0.8$ and up to $T=10^{6}$ Kelvin). The PIMC results are compared to the Drude model in the imaginary domain, where they match within statistical limits and $\Delta\tilde{G}_{\mu\mu}<2\%$, when the Drude damping coefficient $\Gamma$ is small. That is, the PIMC results for $\tilde{G}_{\mu\mu}$ show no statistically significant effects due (reasonable) simulation parameters, such as finite time-step, periodic images or finite number of particles, nor physical parameters such as finite density and temperature. This corroborates that the Drude spectrum is suitable for describing the linear optical response (i.e. the long-wavelength limit) regardless of the coupling strength. Perhaps some minor methodological, systematic biases would have surfaced with generous statistical sampling, but this does not appear as a worthy goal.
@@
In summary: due to the collective motion of many interacting particles in a flat potential, all results come out consistent, effectively, with a free particle. This physical conclusion may not be overly interesting, but it firmly establishes that the numerical method, including the complete post-processing and analysis framework, is valid and robust. To this end, we have demonstrated special sampling approaches (see Sec. II.4), to allow accurate simulation even with a low particle count. These practices form a baseline for numerous further developments:
i) The Fermion spectrum. The treatment of signed permutations should be considered in the future, to unfold the true electron gas behaviour, including complexities in the statistical performance, sampling acceleration and path winding.
@@
ii) Confined spectrum. Treating interacting particles in environments less regular than a flat potential (e.g. weak confinements, optical cavities or ionic plasma) should introduce non-trivial thermal and quantum effects for example with respect to the Drude–Lorentz model.
iii) Inverse Laplace transform. Special methods to invert Eqs. (16) or (17) could be studied to migrate possible quantum corrections into effective models. The MaxEnt method is not the best suited for the broad, Drude-like spectra [23, 41] of this work, but in the presence of resonance peaks, it could be.
@@
iv) Higher-order response properties. The response functions beyond the dipole order could be more sensitive to thermal quantum many-body effects. Their static and dynamic variants, such as $\chi^{(3)}$ of the Drude-like materials, could inform effective models [5] and applications, such as ENZ materials [42].

### v2

@@ IV Results
To make physical benchmarks, there is obvious systematic bias related to the quantum statistics, i.e., the lack of Fermion exchange in the simulation. Otherwise, the Coulomb pair-density matrix is numerically exact and there are no nodal constraints. Remaining bias ensues from using a finite time-step $\Delta\tau$ to treat higher-body correlations (beyond pair approximation) and periodic images. As the higher-body correlations intensify with density, we assume for this work a choice of the time-step according to $\Delta\tau=ar_{s}^{3}$, where $a=0.001$ Ha-1. All these effects are elaborated and further discussed in the SI.
@@
The finite-size effects of the dipole polarizability are more topical: The scaling of $\mathcal{\tilde{G}}(\tau)$ and $\mathcal{\tilde{G}}(i\omega_{n})$ versus $N$ is plotted in Fig. 4 for $r_{c}k_{c}=15.0$ at $r_{s}=4.0$ and $\Theta=0.2$. However, the finite-size effects of $\mathcal{\tilde{G}}$ due to $N$ between PIMC versus Drude are apparently negligible at this density. Nevertheless, we will use $N=16$ for the remainder of this work. The effects due to finite time-step and the Ewald cutoff parameter $r_{c}k_{c}$ are also statistically insignificant and only further discussed in the SI.
@@
Figure 5: On the left, $\mathcal{\tilde{G}}(\tau)$ and $\Delta\mathcal{\tilde{G}}(\tau)$ with variable $\Theta$ at $r_{s}=4$. On the right, $\mathcal{\tilde{G}}(i\omega_{n})$ and $\Delta\mathcal{\tilde{G}}(i\omega_{n})$ with variable $\Theta$ at $r_{s}=4$. Solid lines represent the absolute data and dotted lines the statistical uncertainties.
@@
Figure 6: On the left, a comparison of $\mathcal{\tilde{G}}(\tau)$ versus $\tau$ from PIMC (solid black line) and the Drude model with variable values for $\Gamma$. On the right, the rescaled goodness measure $\eta$ is plotted against $\Gamma$ based on variables temperatures and $r_{s}=4$. Dotted lines are linear fits to the data.
The absence of finite-size effects in $\mathcal{\tilde{G}}$ is attributed to the lack of many-body effects, overall, in the chosen parameter ranges. This is in stark contrast with the energies, indicating that the dipole polarizability is indeed much less sensitive to the dynamic correlation. In Fig. 5 we plot $\mathcal{\tilde{G}}(\tau)$ and $\mathcal{\tilde{G}}(i\omega_{n})$ versus $\Theta$ with $r_{s}=4$ and $N=16$. The plotting axes of $\mathcal{\tilde{G}}(\tau)$ are rescaled to $\tau/\beta$, emphasizing that the absolute values of $\mathcal{\tilde{G}}(\tau=\beta/2)$ decrease with the inverse temperature. Consequently, the Matsubara integral $\mathcal{\tilde{G}}(i\omega_{n}=0)$ goes down with $\beta$, but otherwise the Matsubara series $\mathcal{\tilde{G}}(i\omega_{n}>0)$ follow the same curve, only with different spacings. We present values of $\eta$ from Eq. (46) in Table 1 for various densities and temperatures. The PIMC results match the Drude model in all cases within at least $1\%$ accuracy. On the positive side, this means that the PIMC method is robust and accurate, albeit redundant, on these parameter ranges relevant for simulating metal structures. Again, the Drude response is also valid and unshaken by the account of Coulomb interactions, despite their effects on the energies. Yet, toward higher $r_{s}$ and lower $\Theta$ the value of $\eta$ is statistically different from 0, if only modestly, indicating that $\mathcal{\tilde{G}}(\tau)$ from PIMC is so much larger that $\mathcal{\tilde{G}^{D}}(\tau)$. Effects of the dynamic correlation are expected to intensify toward Wigner crystallization, only expected at much lower densities with $r_{s}\sim 10^{2}$, and lower temperatures [42].
In Figure 6 we discuss how the match depends on the Drude relaxation parameter $\Gamma$. When $\Gamma$ is increased, the Drude correlation function $\mathcal{\tilde{G}}(\tau)$ loses magnitude. The phenomenological explanation is that the scattering rate of the charge carriers from ions increases with $\Gamma$, and thus, the spatial diffusion slows down. The PIMC simulation of free particles has no scattering, and thus, the correspondence is the best with low $\Gamma$. This is presented in the lower panel of Fig. 6, where linear fits approach $\eta\rightarrow 0$ when $\Gamma\rightarrow 0$. The relative error $\eta$ scales linearly with $\Gamma$ and $\beta$, and so, curves representing $\eta/\beta$ are overlaid within statistical uncertainties. For example, observed scattering constants for copper, silver and gold are, respectively, $3.5\times 10^{-3}$, $7.8\times 10^{-4}$ and $2.6\times 10^{-3}$ [4] in atomic units.
With this consideration we want to suggest a future possibility of estimating $\Gamma$ from PIMC simulations, provided a reasonable but yet unknown sampling method of the scattering process. As seen from all the presented results, such a simulation could be easily validated, while providing first-principles insights to the Drude scattering. Furthermore, using a finite-$\Gamma$ simulation such as that could also lend itself for studying non-linear optical phenomena of Drude-like materials on a quantum mechanical level. On the other hand, since finite $\Gamma$ and correlation effects anticipated at high $r_{s}$ both increase $\eta$, they cannot compensate each other. Therefore, it remains to be established, how the many-body effects should be incorporated into an effective and predictive first-principles Drude model.
$r_{s}$
$\Theta$ 2.0 4.0 8.0 12.0
0.1 0.2(0.7) 0.2(0.5) 0.9(0.5) 0.6(0.7)
0.2 -0.0(0.5) -0.1(0.3) 0.4(0.2) 0.4(0.3)
0.4 0.0(0.2) -0.1(0.2) 0.13(11) 0.3(0.2)
0.8 0.1(0.2) -0.1(0.2) 0.12(11) 0.0(0.2)
Table 1: The match parameter $\eta$ with variable $r_{s}$ and $\Theta$ using $N=16$, $r_{c}k_{c}=15$ and $\Gamma=0.0001$.
@@ V Summary and outlook
We have investigated a PIMC estimator for the dipole polarizability of periodic quantum gases and plasmas in the long-wavelength limit, i.e., the optical regime. We have bench marked the estimator with quantum Coulomb plasma in a range of finite densities ($r_{s}=2\ldots 12$ a0) and finite temperatures ($\Theta=0.1\ldots 0.8$ and up to $T=10^{6}$ Kelvin). The PIMC results are compared to the Drude model in the imaginary domain, where they match within statistical limits and $\Delta\mathcal{\tilde{G}}<1\%$, when the Drude $\Gamma$ parameter is small. That is, the PIMC results for $\mathcal{\tilde{G}}$ show no statistically significant effects due (reasonable) simulation parameters, such as finite time-step, periodic images or finite number of particles. Perhaps some minor methodological biases would have surfaced with generous statistical sampling, but this does not appear as a worthy goal.
The physical parameters, finite density and temperature, show almost no effects due to many-body correlation in the chosen parameter ranges. Hints of correlation effects make exceptions at lower densities $r_{s}\geq 8$, i.e., higher coupling constants. On the one hand, this corroborates that the Drude spectrum is effective in describing free charge carriers at room temperatures and metallic densities. On the other hand, revisiting higher coupling constants with proper quantum statistics makes for an intriguing benchmarking prospect.
@@
Even in the absence of many-body effects, it is somehow fascinating to see reduction of the collection of interacting particles into an effectively free particle and a historical, phenomenological model. While this is theoretically anticipated, the numerical side conceals many hurdles which are now firmly established. To this end, we have demonstrated special post-processing, analyses and sampling approaches (see Sec. II.4) to allow accurate simulation even with a low particle count. These practices form a baseline for numerous further developments:
$T$ (K) $\Gamma$ $\eta$ (%)
1000.0 0.0001 1.1(1.1)
0.001 4.3(1.2)
0.01 35(2)
0.1 264(5)
10000.0 0.0001 -0.1(0.3)
0.001 0.2(0.3)
0.01 3.5(0.3)
0.1 33.9(0.3)
100000.0 0.0001 0.06(10)
0.001 0.04(10)
0.01 0.36(10)
0.1 3.62(11)
1000000.0 0.0001 0.07(14)
0.001 0.02(14)
0.01 0.05(14)
0.1 0.39(14)
Table 2: Model matching parameter $\eta$ with statistical uncertainties versus Drude $\Gamma$ parameter at variable fixed temperatures.
i) The Fermion spectrum. The physically accurate treatment of signed permutations should be considered in the future. While this presents challenges with the statistical performance, sampling acceleration and path winding, it also provides an accurate benchmark of quantum many-body corrections in low temperatures and densities.
ii) Inversion to real domain. Where applicable, migration of numerical many-body correction from the integral transformations such as Eqs. (18) or (20) is a formidable challenge. The maximum entropy approaches [27, 25, 21] have been developed, but they work better for sharp Lorentzian peaks than the broad Drude-like spectrum [24, 43]. Finding analytical models and special approaches remains an open challenge.
@@
iii) Confined spectrum. Treating interacting particles in environments less regular than a flat potential (e.g. weak confinements, optical cavities or ionic plasma) should introduce non-trivial thermal and quantum effects for example with respect to the Drude–Lorentz model, characterized with resonance frequencies.
iv) Adsorption on surfaces. Dispersion interactions (e.g. van der Waals and Casimir forces) between atoms and surfaces can be expressed with the instantaneous multipole fluctuations, i.t., polarizabilities in the Matsubara frequency domain [44]. As the PIMC method treats them fluently [29, 30], we may now also use it to inform effective adsorption models for Drude-like matter, such as HEG [45].
@@
v) Higher-order response properties. The response functions beyond the dipole order could be more sensitive to thermal quantum many-body effects. Their static and dynamic variants, such as $\chi^{(3)}$ of the Drude-like materials, could inform effective models [5] and applications, such as ENZ materials [46].
VI Supplemental information
The Supplemental information contains further benchmark figures of accuracy.
@@ Acknowledgements.
Appendix A Reduction of Lindhard theory to Drude model in the long-wavelength limit
The Lindhard function in 3 dimensions, first mentioned in Sec. I, is given by [11]:
$\chi_{0}(\textbf{q},\omega)=\frac{1}{L^{3}}\sum\frac{f_{\textbf{k}+\textbf{q}}{-f_{\textbf{k}}}}{E_{\textbf{k}+\textbf{q}}-E_{\textbf{k}}-\hbar(\omega+i\delta)},$ (49)
where $\chi_{0}(\textbf{q},\omega)$ is the linear response to an incoming electric field with momentum q and frequency $\omega$, k and $E=\hbar^{2}k^{2}/2m_{e}$ are, respectively, the allowed momenta and associated energies in the electron gas, $m_{e}$ is the electron mass, $f_{\mathbf{k}}$ is the Fermi-Dirac distribution, and $\delta$ is an infinitesimal term associated to adiabatic switching-on. In the case of $G=0$, the response given by Eq. (1) is given by
$\begin{split}\chi(\textbf{q},\omega)=&-v_{q}\chi_{0}(\textbf{q},\omega)\\ =&-\frac{v_{q}}{L^{3}}\sum\frac{f_{\textbf{k}+\textbf{q}}{-f_{\textbf{k}}}}{E_{\textbf{k}+\textbf{q}}-E_{\textbf{k}}-\hbar(\omega+i\delta)},\end{split}$ (50)
where $v_{q}$ is the reciprocal Coulomb potential. It is well known that this susceptibility approaches the Drude model with $\Gamma=0$ by taking the limits of $q\rightarrow 0$ (known as the long-wavelength limit[6, 7]) and $\delta\rightarrow 0$ [47]. The following procedure recovers the Drude model with finite $\Gamma$, as define by Eq. (36), by considering a finite $\delta$.
The numerator of the sum in Eq. 50 can be approximated as
$f_{\textbf{k}+\textbf{q}}-f_{\textbf{k}}\simeq\textbf{q}\cdot\nabla_{\textbf{k}}f_{\textbf{k}},$ (51)
while the denominator can be written as
$\begin{split}E_{\textbf{k}+\textbf{q}}-E_{\textbf{k}}&=\frac{\hbar^{2}}{2m_{e}}(k^{2}+2\textbf{k}\cdot\textbf{q}+q^{2})-\frac{\hbar^{2}k^{2}}{2m_{e}}-\hbar(\omega+i\delta)\\ &\simeq\frac{\hbar^{2}\textbf{k}\cdot\textbf{q}}{m_{e}}-\hbar(\omega+i\delta).\end{split}$ (52)
Substituting these approximations in Eq. 50 leads to:
$\chi(\textbf{q},\omega)\simeq-\frac{v_{q}}{L^{3}}\sum_{\textbf{k}}\frac{\textbf{q}\cdot\nabla_{\textbf{k}}f_{\textbf{k}}}{\frac{\hbar^{2}\textbf{k}\cdot\textbf{q}}{m_{e}}-\hbar(\omega+i\delta).}$ (53)
Taking the factor $\hbar(\omega+i\delta)$ outside of the sum and approximating the denominator with its Taylor expansion results in
[…]
In Eq. 54, the term proportional to $\sum_{\textbf{k}}\nabla_{\textbf{k}}f_{\textbf{k}}$ vanishes and the remaining term can be solved by partial integration [47]. Then, the sum $\sum_{\textbf{k}}f_{\textbf{k}}$ corresponds to the number of electrons $N$. Finally, the fact that $v_{q}=e^{2}/|\textbf{q}|^{2}\varepsilon_{0}$ is used in the last step.
At this point, the dependency on $q$ is dropped and the quantity $\omega^{2}_{p}=(ne^{2}/\varepsilon_{0}m_{e})$ is recognized as the plasma frequency with number density $n=N/L^{3}$. Using these definitions leads to
$\chi(\omega)=-\frac{\omega_{p}^{2}}{(\omega+i\delta)^{2}}.$ (55)
Finally, by defining $\delta=\Gamma/2$ and considering $\delta^{2}\approx 0$ , Eq. 37 is recovered
$\chi^{D}(\omega)=-\frac{\omega_{p}^{2}}{\omega^{2}+i\omega\Gamma}.$ (56)
We point that the fact that both $\delta$ in the Lindhard function and $\Gamma$ in the Drude model are infinitesimally small is the reason why we observe the best match in our results for small $\Gamma$, as seen in Table 2.